# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ E.4 Additional Plug-in estimated methods
Table 6: Average detection delay with 3-dimensional time series for a change at time $t=1000$. Comparison of oracle against pre-change mean known methods with plug-in mean estimate
magnitude density FOCuS0 FOCuS0 (est 500) FOCuS0 (est 250) FOCuS0 (est 100) ocd (ora) ocd (est 500) ocd (est 250) ocd (est 100)
@@
Table 7: Average detection delay with 5-dimensional time series for a change at time $t=1000$. Comparison of oracle against pre-change mean known methods with plug-in mean estimate
magnitude density FOCuS0 FOCuS0 (est 500) FOCuS0 (est 250) FOCuS0 (est 100) ocd (ora) ocd (est 500) ocd (est 250) ocd (est 100)
@@
Table 8: Average detection delay with 100-dimensional time series for a change at time $t=1000$. Comparison of oracle against pre-change mean known methods with plug-in mean estimate.
magnitude density FOCuS0 FOCuS0 (est 500) FOCuS0 (est 250) FOCuS0 (est 100) ocd (ora) ocd (est 500) ocd (est 250) ocd (est 100)
@@
Appendix F Dyadic MdFOCuS : an algorithm with a quasi-linear expected complexity
In this appendix we present an algorithm for online changepoint detection. In essence, as in the MdFOCuS algorithm of the main text, it maintains a set of points that is slightly larger than the set of points on the hull. It does so applying the QuickHull (or any convex hull) algorithm on successive and non-overlapping deterministic chunks of the points. This deterministic choice of chunks is likely sub-optimal but it allows for a precise quantification of the expected complexity of the algorithm under the assumption of Theorem 6 (or Corollary 1) and under the assumption that the convex hull algorithm is at worst quadratic. To be specific we obtain a $\mathcal{O}(n\log(n)^{p+1})$ bound on the expected time complexity.
The algorithm we propose (as one of the main text) considers at least all points on the hull to maximise the likelihood for every new data point. For a fixed $p$ following, Theorem 6 and its corollary the number of points on the hull at step $i$ is in $\Omega((\log(i))^{p}).$ To recover a lower bound on the complexity, we first sum overall $i\leq n$, and then consider only $i\geq n/2$ to recover that the overall number of points we will consider is in $\Omega(n/2(\log(n/2))^{p})$ and thus $\Omega(n(\log(n))^{p}).$ Relative to that our algorithm loses just one $\log(n)$ term, we thus argue our bound is relatively tight.
The assumption on the quadratic complexity is true for QuickHull in dimensions 2 and 3 (so for $p\leq 2$) [48, 63] and using the algorithm of Chazelle [64] it is also true in dimension 4 (so for $p\leq 3$). Still following Chazelle this is no longer true for any dimension strictly larger than $4$ and therefore $p>3$.
@@ Appendix F Dyadic MdFOCuS : an algorithm with a quasi-linear expected complexity
This is true under the conditions of Corollary 1.
@@
We assume that the convex hull algorithm we are using is at worst quadratic in time and assume there exist two positive constants $c_{3}$ and $c_{4}$ such that for $n$ points the time complexity is $Time(n)\leq c_{3}n^{2}+c_{4}$.
We now state our bound on the expectation of the square of the number of points on the hull $V_{n}^{p}$.
@@
Now taking the expectation and noting that under the assumption of Corollary 1 the distribution of $|\mathcal{W}_{1}|$ is the same as the distribution of $|\mathcal{W}_{2}|$ we get
@@
F.1 Dyadic algorithm for MdFOCuS
F.1.1 Set of candidates
@@ 20 $\{1,\ldots 16\}$emptyset $\emptyset$ $\{17,18\}$$\{19\}$
Now note that for any $q_{min}\leq\lfloor\log_{2}(n-1)\rfloor$
@@
F.1.2 Updating the set of candidates
To compute $\mathcal{T}_{n+1}$ from scratch we need to apply the hull algorithm on all $\mathcal{U}_{n+1}^{q}$ for $q\geq q_{min}$. Assuming we already computed $\mathcal{T}_{n}$ computing $\mathcal{T}_{n+1}$ is a bit easier. This is because it is often the case that $\mathcal{U}_{n+1}^{q}=\mathcal{U}_{n}^{q}$. Therefore we directly get the indices of the point on the hull of $\mathcal{U}_{n+1}^{q}$ as $\mathcal{U}_{n}^{q}\cap\mathcal{T}_{n}$ because
@@ F.1.2 Updating the set of candidates
Our goal is to get the point (or indices of these points) on the hull of $P(\tau)$ for $\tau$ in $\mathcal{U}_{n+1}^{q_{n}}$. To exploit the merging bound of (20) Algorithm 2 iteratively apply the hull algorithm on larger and larger chunks of the data trimming the points that are not on the hull. In the while loop, from lines 4 to 9, of Algorithm 2 we can distinguish two cases.
@@
Proceeding in such a way at each step we merge the points on the hull of two successive chunks of size $2^{q-1}$ to recover the point on the hull of a chunk of size $2^{q}$. Figure 14 is a schematic representation of these successive fusions (applications of the hull algorithm).
@@
Under the assumption of Corollary 1 (see Assumption 1) and Assumption 2, the expected time complexity of Algorithm 2 is
[…]
@@
For the smallest scale, $n^{\prime}=2^{q_{min}},$ we have $n/2^{q_{min}}$ chunks and run QuickHull on all points. Using Assumption 2, the worst time per chunk is less than $c_{3}(2^{q_{min}})^{2}+c_{4},$ and summing overall chunks we get $nc_{3}2^{q_{min}}+\frac{nc_{4}}{2^{q_{min}}}$ operations. So the smallest scale is in $\mathcal{O}(n).$
We will now successively consider all larger scales $n^{\prime}=2^{q}$ for $q_{min}<q\leq\lfloor\log_{2}(n)\rfloor.$ Note that the hull of a chunk of size $n^{\prime}=2^{q}$ is obtained by combining the hulls of two smaller chunks of size $2^{q-1}.$ For any such chunk we can thus exploit Lemma 4 and equation (20) with $|\mathcal{W}_{1}|=|\mathcal{W}_{2}|=n^{\prime}/2=2^{q-1}$ and we get
@@
Multiplying by $n/2^{q}$ to consider all chunks of size $2^{q}$ we get that the time to process scale $2^{q}$ is
$\mathbb{E}\left(Time(\text{``Scale }2^{q}\text{''})\right)$ ” ) ) $\leq$ $\frac{n}{2^{q}}\left(4c_{1}c_{3}2^{q-1}(\log 2^{q-1})^{p}+4c_{2}c_{3}2^{q-1}+c_{4}\right)$
$\leq$ $4c_{1}c_{3}\frac{n}{2}(\log 2^{q-1})^{p}+4c_{2}c_{3}\frac{n}{2}+c_{4}\frac{n}{2^{q}}$
@@
So scale $2^{q}$ is in $\mathcal{O}(n\log(n)^{p}).$ Multiplying by $\log_{2}(n)$ to consider all scales we get the desired result. ∎
F.1.3 MdFOCuS with dyadic update
The MdFOCuS Algorithm implementing this update of the set of candidates is presented in Algorithm 3. Combining Lemma 5 and 2 we get an overall expected complexity in $\mathcal{O}(n\log(n)^{p+1})$ complexity.
Algorithm 3 MdFOCuS(0) Dyadic algorithm
@@ F.1.3 MdFOCuS with dyadic update
3:Input 3: $q_{min}$ min power 2 scale at which we run QuickHull
4:Output: stopping time $n$ and maximum likelihood change: $\hat{\tau}(.)$ or $\hat{\tau}(\eta_{1})$
5:$logLikelihoodRatio\leftarrow-\infty,\quad\mathcal{T}\leftarrow\emptyset,\quad n\leftarrow 0$
6:while ($logLikelihoodRatio<thrs$do
7:     $n\leftarrow n+1$
8:     $\mathcal{T}\leftarrow\mathcal{T}\cup\{n-1\}$
9:     $maxLogLikelihood\leftarrow\max_{\tau\in\mathcal{T}}\{\max_{\eta_{1},\eta_{2}}\ell_{\tau,n}(\eta_{1},\eta_{2})\}$triangleright $\eta_{1}$ is fixed in MdFOCuS0
10:     $logLikelihoodRatio\leftarrow maxLogLikelihood-\max_{\eta_{1},\eta_{2}}\ell_{n,n}(\eta_{1},\eta_{2})$triangleright $\eta_{1}$ is fixed in MdFOCuS0
11:     $q\leftarrow q_{min}$
12:     while $remainder(\frac{n-1}{2^{q}})=0$  do
13:         $\mathcal{T}^{\prime}\leftarrow\{\tau\in\mathcal{T}|\tau>n-2^{q}\}$
14:         $\mathcal{T}^{\prime\prime}\leftarrow\textsc{Quickhull}(\{P(\tau)\}_{\tau\in\mathcal{T}^{\prime}})$
[…]
16:         $q\leftarrow q+1$
17:     end while
18:end while
19:return $n$ and $logLikelihoodRatio$
We implemented Algorithm 3 also for a change in the mean of multi-dimensional Gaussian signal (see Table 1) in R/C++ using the qhull library (see link in the Supplementary Material).
We evaluate the empirical run times of Algorithm 3 with various values of $q_{min}$ using the profiles of Supplementary Material E.1. In Figure 16 we report the average run time as a function of the dyadic parameter $q_{min}$ for MdFOCuS0 with dyadic update (that is for a known value of $\eta_{1}$) and values of $q_{min}$ in ${3,\dots,12}$. We observe a minimum run time for $q_{min}$ in $6$ for $p=1$, in $7$ for $p=2$ and in $8$ for $p=3$. As a consequence in the rest of the simulations we always set $q_{min}$ to $p+5$.
Using the simulation framework of Section 4.3 we evaluated the run times of MdFOCuS0 with dyadic update for dimension $p=1$ to $p=3$ and compared it with run times of MdFOCuS. The average run times are presented in Figure 16. Dyadic MdFOCuS0 is slower than MdFOCuS0, about 2 times slower.
Figure 15: Run time of MdFOCuS0 with dyadic update as a function of $q_{min}$ for $p$-variate time series with $n=10^{5}$ data points in dimension $p=1,2$ and $3$. We simulated $100$ i.i.d. Gaussian data $\mathcal{N}_{p}(0,I_{p})$ and report the average run time.
Figure 16: Run times in seconds of MdFOCuS0 (dotted) and MdFOCuS0 with dyadic update (full) both with $q_{min}=5+p$ in dimension $p=1,2$ and $3$ using time series $x_{1:n}$ (without change, $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$ i.i.d.). Run times are averaged over $100$ data sets. We considered a maximum running time of $10$ minutes for the algorithms. That is why we do not report run times for $p=3$ for large signals.
Appendix G An approximation for large $p$
Based on Theorem 6 and its corollary at step $n$ we expect Algorithm 1 to store $\mathcal{O}(\log^{p}(n))$ candidates. Considering the calculation of the full and sparse maximum likelihood in lines 9 and 10 should thus be in $\mathcal{O}(p\log^{p}(n)).$ The power $\log(n)^{p}$ is problematic making the algorithm impractical for $p$ larger than $6$, as exemplified in Figure 4. Here we propose a heuristic solution to this problem. Our idea is to get a subset of the point on the hull considering projections on at most $\tilde{p}$ variates.
More formerly for any ordered subset $S=\{S_{1},\ldots,S_{|S|}\}$ of $\{1,\ldots,p\}$, with $|S|$ the size of $S$, we define $x_{t}^{[S]}$ the vector in $\mathbb{R}^{|S|}$ such that the $i-th$ coordinate of $x_{t}^{[S]}$ is the $S_{i}$-coordinate of $x_{t}$. We then define $P^{[S]}(\tau)$ accordingly:
$P^{[S]}(\tau)=\left(\tau,\sum_{t=1}^{\tau}x_{t}^{[S]}\right),$
$\cup_{S\in\mathcal{S}}\mathcal{T}^{[S]}_{n}\subseteq\mathcal{T}_{n}.$
Based on Theorem 6 and its corollary at step $n$ $E(|\mathcal{T}^{[S]}_{n}|)=\mathcal{O}(\log(n)^{|S|})$. The quality of the approximation is thus be governed by $|\mathcal{S}|$ and $\max_{S\in\mathcal{S}}(|S|).$ The resulting procedure is formally described in Algorithm 4.
In our simulations as a minimalist implementation of this approximation we considered $\tilde{p}=2$ and subsets of the form $\{\tilde{p}j+1,\ldots,\tilde{p}j+\tilde{p}\}$ for $j\leq p/\tilde{p}.$ We have $p/\tilde{p}$ such sets and under the condition of Theorem 6 we thus expect to store $\mathcal{O}(p\log^{\tilde{p}}(n))$ candidate changepoints rather than $\mathcal{O}(\log^{p}(n))$ with the exact Algorithm 1. For $p=100$ and $n=10000$, our approximation with $\tilde{p}=2$, called MdFOCuS 2d, implemented in R runs on average in less than 100 seconds. More generally it can be seen in Figure 5 that run times of our R implementation of MdFOCuS, and our 2d approximation for larger $p$ are comparable to ocd [38].
Algorithm 4 MdFOCuS(0) $\tilde{p}$d approximation
1:Input 1: $p$-variate independent time series $\{x_{t}\}_{t=1,2,\dots}$
2:Input 2: threshold $thrs$ and $\eta_{1}$ if known pre-change parameter
3:Input 3: $\mathcal{S}$ a subset of all subsets of $\{1,\ldots,p\}$ of size $\tilde{p}$
@@ Appendix G An approximation for large $p$
10:     For all $S$$\mathcal{T}^{[S]}\leftarrow\mathcal{T}^{[S]}\cup\{n-1\}$
11:     $maxLogLikelihood\leftarrow\max_{\tau\in\mathcal{T}}\{\max_{\eta_{1},\eta_{2}}\ell_{\tau,n}(\eta_{1},\eta_{2})\}$triangleright $\eta_{1}$ is fixed in MdFOCuS0
12:     $logLikelihoodRatio\leftarrow maxLogLikelihood-\max_{\eta_{1},\eta_{2}}\ell_{n,n}(\eta_{1},\eta_{2})$triangleright $\eta_{1}$ is fixed in MdFOCuS0
13:     for all $S\in\mathcal{S}$ do
14:         if ($|\mathcal{T}^{[S]}|>maxSize^{[S]}$then
15:              $\mathcal{\mathcal{missing}}{T}^{[S]}\leftarrow\textsc{Quickhull}(\{P^{[S]}(\tau)\}_{\tau\in\mathcal{T}^{[S]}})$
16:              $maxSize^{[S]}\leftarrow\lfloor\alpha|\mathcal{T}^{[S]}|+\beta\rfloor$
17:         end if
18:     end for
19:end while
20:return $n$ and $logLikelihoodRatio$
@@
Appendix H On the probability of false alarms, average-run-length and average detection delay in the multivariate Gaussian case

### v3

@@ F.4 Additional Plug-in estimated methods
Table 6: Average detection delay with 3-dimensional time series for a change at time $t=1000$. Comparison of pre-change mean known methods with plug-in oracle value (ora) against plug-in mean estimate for differing training data sizes (500, 250 and 100).
magnitude sparsity FOCuS (ora) FOCuS (est 500) FOCuS (est 250) FOCuS (est 100) ocd (ora) ocd (est 500) ocd (est 250) ocd (est 100)
@@
Table 7: Average detection delay with 5-dimensional time series for a change at time $t=1000$. Comparison of pre-change mean known methods with plug-in oracle value (ora) against plug-in mean estimate for differing training data sizes (500, 250 and 100).
magnitude sparsity FOCuS (ora) FOCuS (est 500) FOCuS (est 250) FOCuS (est 100) ocd (ora) ocd (est 500) ocd (est 250) ocd (est 100)
@@
Table 8: Average detection delay with 100-dimensional time series for a change at time $t=1000$. Comparison of pre-change mean known methods with plug-in oracle value (ora) against plug-in mean estimate for differing training data sizes (500, 250 and 100).
magnitude sparsity FOCuS (ora) FOCuS (est 500) FOCuS (est 250) FOCuS (est 100) ocd (ora) ocd (est 500) ocd (est 250) ocd (est 100)
@@
Appendix G Dyadic MdFOCuS : an algorithm with a quasi-linear expected complexity
In this appendix we present an algorithm for online changepoint detection. In essence, as in the MdFOCuS algorithm of the main text, it maintains a set of points that is slightly larger than the set of points on the hull. It does so applying the QuickHull (or any convex hull) algorithm on successive and non-overlapping deterministic chunks of the points. This deterministic choice of chunks is likely sub-optimal but it allows for a precise quantification of the expected complexity of the algorithm under the assumption of Theorem 6 and under the assumption that the convex hull algorithm is at worst quadratic. To be specific we obtain a $\mathcal{O}(n\log(n)^{p+1})$ bound on the expected time complexity.
@@ Appendix G Dyadic MdFOCuS : an algorithm with a quasi-linear expected complexity
This is true under the conditions of Theorem 6 and Corollary 1.
Remember we define $Time(n)$ to be the worst-case time complexity of finding the convex hull of $n$ points.
@@
For our convex hull algorithm we are using there exist two positive constants $c_{3}$ and $c_{4}$ such that $Time(n)\leq c_{3}n^{2}+c_{4}$.
The assumption on the quadratic complexity is true for QuickHull for $p\leq 2$ Barber et al. [1996] and using the algorithm of Chazelle [1993] it is true for $p\leq 3$. We now state our bound on the expectation of the square of the number of points on the hull $V_{n}^{p}$.
@@
Now taking the expectation with respect to the number of points in $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$, and noting that under the assumption of Corollary 1 the distribution of $|\mathcal{W}_{1}|$ is the same as the distribution of $|\mathcal{W}_{2}|$ we get
@@
Now, before presenting the dyadic algorithm, we explain the set of candidates it considers and how this set is updated.
G.1 Dyadic algorithm: Set of candidates
@@ 20 $\{1,\ldots 16\}$emptyset $\emptyset$ $\{17,18\}$$\{19\}$
For any $q_{min}\leq\lfloor\log_{2}(n-1)\rfloor$
@@
G.2 Dyadic algorithm: Updating the set of candidates
To compute $\mathcal{T}_{n+1}$ from scratch we need to apply the hull algorithm on all $\mathcal{U}_{n+1}^{q}$ for $q\geq q_{min}$. Assuming we have already computed $\mathcal{T}_{n}$ computing $\mathcal{T}_{n+1}$ is a bit easier. This is because it is often the case that $\mathcal{U}_{n+1}^{q}=\mathcal{U}_{n}^{q}$. Therefore we directly get the indices of the point on the hull of $\mathcal{U}_{n+1}^{q}$ as $\mathcal{U}_{n}^{q}\cap\mathcal{T}_{n}$ because
@@ G.2 Dyadic algorithm: Updating the set of candidates
Our goal is to get the point (or indices of these points) on the hull of $P(\tau)$ for $\tau$ in $\mathcal{U}_{n+1}^{q_{n}}$. To exploit the merging bound of (17) Algorithm 2 iteratively applies the hull algorithm on larger and larger chunks of the data trimming the points that are not on the hull. In the while loop, from lines 4 to 9, of Algorithm 2 we can distinguish two cases.
@@
Proceeding in such a way at each step we merge the points on the hull of two successive chunks of size $2^{q-1}$ to recover the points on the hull of a chunk of size $2^{q}$. Figure 14 is a schematic representation of these successive fusions (applications of the hull algorithm).
@@
Under Assumptions 1 and 2, the expected time complexity of Algorithm 2 is $\mathcal{O}\left(n\log(n)^{p+1}\right)$.
@@
For the smallest scale, $n^{\prime}=2^{q_{min}},$ we have $n/2^{q_{min}}$ chunks and run QuickHull on all points. Using Assumption 2, the worst time per chunk is less than $c_{3}(2^{q_{min}})^{2}+c_{4},$ and summing overall chunks we get $nc_{3}2^{q_{min}}+\frac{nc_{4}}{2^{q_{min}}}$ operations. So at the smallest, scale the computational cost is $\mathcal{O}(n)$.
We will now successively consider all larger scales $n^{\prime}=2^{q}$ for $q_{min}<q\leq\lfloor\log_{2}(n)\rfloor$. The hull of a chunk of size $n^{\prime}=2^{q}$ is obtained by combining the hulls of two smaller chunks of size $2^{q-1}$. For any such chunk, we can thus exploit Lemma 4 and (17) with $|\mathcal{W}_{1}|=|\mathcal{W}_{2}|=n^{\prime}/2=2^{q-1}$ and we get
@@
Multiplying by $n/2^{q}$ to consider all chunks of size $2^{q}$, we get that the time to process scale $2^{q}$ is
$\frac{n}{2^{q}}\left(4c_{1}c_{3}2^{q-1}(\log 2^{q-1})^{p}+4c_{2}c_{3}2^{q-1}+c_{4}\right)$$\leq$ $4c_{1}c_{3}\frac{n}{2}(\log 2^{q-1})^{p}+4c_{2}c_{3}\frac{n}{2}+c_{4}\frac{n}{2^{q}}$
@@
So the computational cost at scale $2^{q}$ is $\mathcal{O}(n\log(n)^{p})$. Multiplying by $\log_{2}(n)$ to consider all scales, we get the desired result. ∎
G.3 MdFOCuS with dyadic update
The MdFOCuS Algorithm implementing this update of the set of candidates is presented in Algorithm 3. Combining Lemma 5 and 1 we get an overall complexity that is $\mathcal{O}(n\log(n)^{p+1})$.
Algorithm 3 MdFOCuS Dyadic algorithm
1:Input 1: $\{x_{t}\}_{t=1,2,\dots}$ $p$-variate independent time series; $thrs$ threshold; $\eta_{1}$ pre-change parameter (if known); $q_{min}$ min power 2 scale at which we run QuickHull
2:Outputs: stopping time $n$ and maximum likelihood change: $\hat{\tau}(.)$ or $\hat{\tau}(\eta_{1})$
3:$GLR\leftarrow-\infty,\quad\mathcal{T}\leftarrow\emptyset,\quad n\leftarrow 1$
4:while ($GLR<thrs$do
5:     $n\leftarrow n+1$
6:     $\mathcal{T}\leftarrow\mathcal{T}\cup\{n-1\}$
7:     $\hat{m}\leftarrow\max_{\tau\in\mathcal{T}}\{\max_{\eta_{1},\eta_{2}}\ell_{\tau,n}(\eta_{1},\eta_{2})\}$triangleright $\eta_{1}$ is fixed if pre-change known
8:     $GLR\leftarrow\hat{m}-\max_{\eta_{1},\eta_{2}}\ell_{n,n}(\eta_{1},\eta_{2})$triangleright $\eta_{1}$ is fixed if pre-change known
9:     $q\leftarrow q_{min}$
10:     while $\textsc{remainder}(\frac{n-1}{2^{q}})=0$  do
11:         $\mathcal{T}^{\prime}\leftarrow\{\tau\in\mathcal{T}|\tau>n-2^{q}\}$
12:         $\mathcal{T}^{\prime\prime}\leftarrow\textsc{Quickhull}(\{P(\tau)\}_{\tau\in\mathcal{T}^{\prime}})$
[…]
14:         $q\leftarrow q+1$
15:     end while
16:end while
17:return $n$ and $GLR$
We implemented Algorithm 3 for a change in the mean of multi-dimensional Gaussian signal in R/C++ using the qhull library. We evaluate the empirical run times of Algorithm 3 with various values of $q_{min}$ using the profiles of Supplementary Material F.2. In Figure 16 we report the average run time as a function of the dyadic parameter $q_{min}$ for MdFOCuS for a known value of $\eta_{1}$ with the dyadic update and values of $q_{min}$ in ${3,\dots,12}$. We observe a minimum run time for $q_{min}=6$ for $p=1$, $q_{min}=7$ for $p=2$ and $q_{min}=8$ for $p=3$. As a consequence in the rest of the simulations we always set $q_{min}$ to $p+5$.
Using the simulation framework of Section 1 we evaluated the run times of MdFOCuS with dyadic update for dimension $p=1$ to $p=3$ and compared it with run times of MdFOCuS. The average run times are presented in Figure 16. Dyadic MdFOCuS is slower than MdFOCuS, by about a factor of 2.
Figure 15: Run time of MdFOCuS with dyadic update as a function of $q_{min}$ for $p$-variate time series with $n=10^{5}$ data points in dimension $p=1,2$ and $3$. We simulated $100$ i.i.d. Gaussian data $\mathcal{N}_{p}(0,I_{p})$ and report the average run time.
Figure 16: Run times in seconds of MdFOCuS (dotted) and MdFOCuS with dyadic update (full) both with $q_{min}=5+p$ in dimension $p=1,2$ and $3$ using time series $x_{1:n}$ (without change, $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$ i.i.d.). Run times are averaged over $100$ data sets. We considered a maximum running time of $10$ minutes for the algorithms. That is why we do not report run times for $p=3$ for large signals.
Appendix H An approximation for large $p$
Based on Theorem 6 and its corollary at step $n$ we expect Algorithm 1 to store $\mathcal{O}(\log^{p}(n))$ candidates. Considering the calculation of the full and sparse maximum likelihood in lines 7 and 8 should thus have a complexity of $\mathcal{O}(p\log^{p}(n)).$ The power $\log(n)^{p}$ is problematic making the algorithm impractical for $p$ larger than $6$, as exemplified in Figure 3.
Here we propose a heuristic solution to this problem. Our idea is to get a subset of the point on the hull considering projections on at most $\tilde{p}$ variates.
For any ordered subset $S=\{S_{1},\ldots,S_{|S|}\}$ of $\{1,\ldots,p\}$, with $|S|$ the size of $S$, we define $x_{t}^{[S]}$ the vector in $\mathbb{R}^{|S|}$ such that the $i-th$ coordinate of $x_{t}^{[S]}$ is the $S_{i}$-coordinate of $x_{t}$. We then define $P^{[S]}(\tau)$ accordingly:
$P^{[S]}(\tau)=\left(\tau,\sum_{t=1}^{\tau}x_{t}^{[S]}\right).$
$\cup_{S\in\mathcal{S}}\mathcal{T}^{[S]}_{n}\subseteq\mathcal{T}_{n}.$
In our simulations as a minimalist implementation of this approximation we considered $\tilde{p}=2$ and subsets of the form $\{\tilde{p}j+1,\ldots,\tilde{p}j+\tilde{p}\}$ for $j\leq p/\tilde{p}.$ We have $p/\tilde{p}$ such sets and under the condition of Theorem 6 we thus expect to store $\mathcal{O}(p\log^{\tilde{p}}(n))$ candidate changepoints rather than $\mathcal{O}(\log^{p}(n))$ with the exact Algorithm 1. For $p=100$ and $n=10000$, our approximation with $\tilde{p}=2$, implemented in R runs on average in less than 100 seconds. More generally it can be seen in Figure 4 that run times of our R implementation of MdFOCuS, and our 2d approximation for larger $p$ are comparable to ocd Chen et al. [2022].
Algorithm 4 MdFOCuS $\tilde{p}$d approximation
@@ Appendix H An approximation for large $p$
3:Input 3: $\mathcal{S}$ a subset of all subsets of $\{1,\ldots,p\}$, with for all $S\in\mathcal{S}$ : $|S|\leq\tilde{p}$
@@
10:     for all $S\in\mathcal{S}$ do
11:         $\mathcal{T}^{[S]}\leftarrow\mathcal{T}^{[S]}\cup\{n-1\}$
12:     end for
13:     $\hat{m}\leftarrow\max_{\tau\in\cup_{S\in\mathcal{S}}\mathcal{T}^{[S]}}\{\max_{\eta_{1},\eta_{2}}\ell_{\tau,n}(\eta_{1},\eta_{2})\}$triangleright $\eta_{1}$ is fixed if pre-change known
14:     $GLR\leftarrow\hat{m}-\max_{\eta_{1},\eta_{2}}\ell_{n,n}(\eta_{1},\eta_{2})$triangleright $\eta_{1}$ is fixed if pre-change known
15:     for all $S\in\mathcal{S}$ do
16:         if ($|\mathcal{T}^{[S]}|>maxSize^{[S]}$then
17:              $\mathcal{T}^{[S]}\leftarrow\textsc{Quickhull}(\{P^{[S]}(\tau)\}_{\tau\in\mathcal{T}^{[S]}})$
18:              $maxSize^{[S]}\leftarrow\lfloor\alpha|\mathcal{T}^{[S]}|+\beta\rfloor$
19:         end if
20:     end for
21:end while
22:return $n$ and $GLR$
@@
Appendix I On the probability of false alarms, average-run-length and average detection delay in the multivariate Gaussian case
I.1 Definitions and problem statement